OpenStudy (anonymous):

There were 150 freshman and sophomores at the high school musical Reese and her friends attended. Freshman tickets were $2.00 each and sophomore tickets were $3.00 each. The receipts for freshman and sophomores tickets totaled $410. How many sophomores bought tickets?

OpenStudy (anonymous):

The same again solve the 2 sets of equations F + S = 150, therefore F = 150 - S The second equation is $2 x F + $3 x S = 410 or 2F + 3S = 410 We know F = 150 - S Therefore 2(150 - S) + 3S = 410 300 -2S + 3S = 410 S = 110 Substitute this back into F = 150 - S and we find that F = 40 Check 40 +110 = 150 tickets - OK 40 x $2 = $80 110 x $3 = $330 $80 + $330 = $410 = O